# WF-VAE: Enhancing Video VAE by Wavelet-Driven Energy Flow for Latent Video Diffusion Model

### v2

@@ 4.2 Comparison With Baseline Models
We compare WF-VAE with baseline models in three key aspects: computational efficiency, reconstruction performance, and performance in diffusion-based generation. To ensure fairness in comparing metrics and model efficiency, we disable block-wise inference strategies across all VAEs.
Computational Efficiency. The computational efficiency evaluations are conducted using an H100 GPU with float32 precision. Performance evaluations are performed at 33 frames across multiple input resolutions. Allegro VAE is evaluated using 32-frame videos to maintain consistency due to its non-causal convolution architecture. For a fair comparison, all benchmark VAEs do not employ block-wise inference strategies and are directly inferred. As shown in Fig. 4, WF-VAE demonstrates superior inference performance compared to other VAEs. For instance, WF-VAE-L requires 7170 MB of memory for encoding a video at 512×512 resolution, whereas OD-VAE demands approximately 31944 MB (445% higher). Similarly, CogVideoX consumes around 35849.33 MB (499% higher), and Allegro VAE requires 55664 MB (776% higher). These results highlight WF-VAE’s significant advantages in large-scale training and data processing. In terms of encoding speed, WF-VAE-L achieves an encoding time of 0.0513 seconds, while OD-VAE, CogVideoX, and Allegro VAE exhibit encoding times of 0.0945 seconds, 0.1810 seconds, and 0.3731 seconds, respectively, which are approximately 184%, 352%, and 727% slower. From Fig. 4, it can also be observed that WF-VAE has significant advantages in decoding computational efficiency.
@@
Video Reconstruction Performance. We present a quantitative comparison of reconstruction performance between WF-VAE and baseline models in Tab. 1 and qualitative reconstruction results in Fig. 5. As shown, despite having the lowest computational cost, WF-VAE-S outperforms popular open-source video VAEs such as OD-VAE [6] and Open-Sora VAE [45] on both datasets. When increasing the model complexity, WF-VAE-L competes well with Allegro [46], outperforming it in PSNR, LPIPS and FVD, but slightly lagging in SSIM. However, WF-VAE-L is significantly more computationally efficient than Allegro. Additionally, we compare WF-VAE-L with CogVideoX [39] using 16 latent channels. Except for a slightly lower SSIM on the Webvid-10M, WF-VAE-L outperform CogVideoX across all other metrics. These results indicate that WF-VAE achieves competitive reconstruction performance compared to state-of-the-art open-source video VAEs, while substantially improving computational efficiency.
Video Generation Evaluation.Fig. 6 and Tab. 2 qualitatively and quantitatively demonstrate the video generation results of the diffusion model using WF-VAE. WF-VAE achieves the best performance in terms of FVD and IS metrics. For instance, in the SkyTimelapse dataset, among models with 4 latent channels, WF-VAE-S achieves the best FVD score, which is 10.23 lower than WF-VAE-L and 27.35 lower than OD-VAE. For models with 16 latent channels, WF-VAE-L’s FVD score is 0.51 lower than CogVideoX’s. This might be because the higher dimensionality of the latent space makes convergence more difficult, resulting in slightly inferior performance compared to WF-VAE with 4 latent channels under the same training steps.
@@ 4.3 Ablation Study
Increasing the latent dimension. Recent works [10, 8] show that the number of latent channel significantly impacts reconstruction quality. We conduct experiments with 4, 8, 16, and 32 latent channels. Fig. 7(a) illustrates that reconstruction performance improves significantly as the number of latent channels increases. However, larger latent channels may increase convergence difficulty in training diffusion model, as evidenced by the results presented in  Tab. 2.
Exploration of WL loss weight $\lambda_{W\!L}$. To ensure structural symmetry in WF-VAE, we introduce WL loss. As shown in Fig. 7(b), the model exhibits a substantial decrease in PSNR performance when $\lambda_{W\!L}=0$. Our experiments demonstrate optimal results for both PSNR and LPIPS metrics when $\lambda_{W\!L}=0.1$. Furthermore, our analysis of different loss functions reveals that utilizing L1 loss produces superior results compared to L2 loss.
@@
Ablation studies of model architecture. First, We evaluate the effectiveness of the energy pathway by examining its impact when removed from layer 3 alone and both layers 2 and 3. This analysis demonstrates the benefits of incorporating low-frequency video energy into the backbone. Second, we investigate the significance of our proposed WL Loss in regularizing the encoder-decoder Third, we examine the impact of modifying the normalization method to layer normalization [1] for Causal Cache. The results of these ablation studies are shown in Tab. 4.
@@ 4.4 Causal Cache
To validate the lossless inference capability of Causal Cache, we compared our approach with existing block-wise inference methods implemented in several open-source LVDMs. OD-VAE [6] and Allegro [46] provide spatio-temporal tiling inference implementations, and CogVideoX [39] implements a temporal caching strategy. As demonstrated in Tab. 5, both tiling strategies and conventional caching methods exhibited performance degradation, while Causal Cache achieves lossless inference with performance metrics identical to direct inference.
@@ 5 Conclusion
Supplementary Material
The supplementary materials include further details as follows:
  • We present additional notations in Sec. 6.
  • We analyze subband energy and entropy of wavelet transform in Sec. 7, which further validates our motivation.
  • We present our training parameters in Sec. 8
  • We present the derivation of the Causal Cache formulation in Sec. 9.
6 Notations
The notations and their descriptions in the paper are shown in Tab. 6.
Notations Descriptions
$WT(\cdot)$ Wavelet transform
$IWT(\cdot)$ Inverse wavelet transform
[…]
$\mathcal{W}^{(l)}$ The set of all subbands within layer $l$
Table 6: Notations symbols and their descriptions.
7 Wavelet Subband Analysis
We analyze the energy and entropy distributions across the subbands obtained after wavelet transform. As illustrated in Fig. 8(b), the energy and entropy of the video are primarily concentrated in the $hhh$ low-frequency subband. This concentration suggests that low-frequency components carry more significant information and necessitate lower compression rates to ensure superior reconstruction performance. This observation further validates the rationale behind our proposed approach.
(a) Visualization of the eight subbands obtained after wavelet transform of the video.
(b) Energy and entropy of each subband.
Figure 8: Visualization of the subbands and their respective energy and entropy.
8 Training Details
The training hyperaparameters are shown in Tab. 7.
Parameter Setting
Stage I - 800k step
Learning Rate 1e-5
Total Batch Size 8
Peceptual(LPIPS) Weight 1.0
WL Loss Weight ($\lambda_{W\!L}$) 0.1
KL Weight ($\lambda_{K\!L}$) 1e-6
Learning Rate 1e-5
Resolution 256$\times$256
Num Frames 25
EMA Decay 0.999
Stage II - 200k step
Num Frames 49
Stage III - 200k step
Peceptual(LPIPS) Weight 0.1
Table 7: Training hyperparameters across three stages.
Figure 9: Illustration of Causal Cache with parameters $k$=3, $s$=2, and chunk size $T_{chunk}$=4.
9 Derivation of Causal Cache
Let us define a convolution with sliding window index $n\in\mathbb{N}_{0}$ and chunk index $m\in\mathbb{N}_{0}$. Given a convolutional stride $s$ and kernel size $k$, as shown in Fig. 9, the starting and ending frame indices for each sliding window are expressed as follows:
$t_{window,start}(n)$ $=ns,$ (8)
$t_{window,end}(n)$ $=t_{window,start}(n)+k-1.$ (9)
For chunk boundaries, we define:
$t_{chunk,end}(m)=k-1+mT_{chunk}$ (10)
where $T_{chunk}$ denotes the chunk size. For a given chunk index $m$, the maximum sliding index $n_{max}(m)$ is determined by the constraint $t_{window,end}(n)\geq t_{chunk,end}(m)$:
$n_{\text{max}}(m)=\left\lfloor\frac{mT_{chunk}}{s}+1\right\rfloor.$ (11)
Consequently, the required cache size $T_{\text{cache}}(m)$ for chunk $m$ is:
$\begin{split}T_{\text{cache}}(m)=t_{\text{chunk,end}}(m)-t_{\text{window,start}}(n_{\text{max}}(m))\\ +1=mT_{\text{chunk}}+k-\left\lfloor\frac{mT_{\text{chunk}}}{s}+1\right\rfloor s\end{split}$ (12)

### v3

@@ 4.2 Comparison With Baseline Models
We compare WF-VAE with baseline models in three key aspects: computational efficiency, reconstruction performance, and diffusion-based generation performance. To ensure fairness in comparing metrics and model efficiency, we disable block-wise inference strategies across all VAEs.
Computational Efficiency. The computational efficiency evaluations are conducted using an H100 GPU with float32 precision. Performance evaluations are performed at 33 frames across multiple input resolutions. Due to non-causal convolution architecture, Allegro VAE is evaluated using 32-frame videos to maintain consistency. All benchmark VAEs use direct inference without block-wise inference strategies for a fair comparison. As shown in Fig. 4, WF-VAE demonstrates superior inference performance compared to other VAEs. For instance, WF-VAE-L requires 7170 MB of memory for encoding a video at 512×512 resolution, whereas OD-VAE demands approximately 31944 MB (445% higher). Similarly, CogVideoX consumes around 35849.33 MB (499% higher), and Allegro VAE requires 55664 MB (776% higher). These results highlight WF-VAE’s significant advantages in large-scale training and data processing. In terms of encoding speed, WF-VAE-L achieves an encoding time of 0.0513 seconds, while OD-VAE, CogVideoX, and Allegro VAE exhibit encoding times of 0.0945 seconds, 0.1810 seconds, and 0.3731 seconds, respectively, which are approximately 184%, 352%, and 727% slower. As illustrated in Fig. 4, WF-VAE demonstrates notable advantages in computational efficiency during the decoding process.
@@
Video Reconstruction Performance. We present a quantitative comparison of reconstruction performance between WF-VAE and baseline models in Tab. 1 and qualitative reconstruction results in Fig. 5. Despite having the lowest computational cost, WF-VAE-S outperforms popular open-source video VAEs such as OD-VAE [6] and Open-Sora VAE [49] on both datasets. When increasing the model complexity, WF-VAE-L competes well with Allegro [50], outperforming it in PSNR, LPIPS and FVD but slightly lagging in SSIM. However, WF-VAE-L is significantly more computationally efficient than Allegro. Additionally, we compare WF-VAE-L with CogVideoX [40] using 16 latent channels. Except for a slightly lower SSIM on the Webvid-10M, WF-VAE-L outperforms CogVideoX across all other metrics. These results indicate that WF-VAE achieves competitive reconstruction performance compared to state-of-the-art open-source video VAEs, while substantially improving computational efficiency.
Video Generation Evaluation.Fig. 6 and Tab. 2 qualitatively and quantitatively demonstrate the video generation results of the diffusion model using WF-VAE. WF-VAE achieves the best performance in terms of FVD and IS metrics. For instance, in the SkyTimelapse dataset, among models with 4 latent channels, WF-VAE-S achieves the best FVD score, 10.23 lower than WF-VAE-L and 27.35 lower than OD-VAE. For models with 16 latent channels, WF-VAE-L’s FVD score is 0.51 lower than CogVideoX’s. This might be because the higher dimensionality of the latent space makes convergence more difficult, resulting in slightly inferior performance compared to WF-VAE with 4 latent channels under the same training steps.
@@ 4.3 Ablation Study
Increasing the latent dimension. Recent works [9, 8] show that the number of latent channels significantly impacts reconstruction quality. We experiment with 4, 8, 16, and 32 latent channels. Fig. 7(a) illustrates that reconstruction performance improves significantly as the number of latent channels increases. However, larger latent channels may increase convergence difficulty in training diffusion model, as evidenced by the results presented in  Tab. 2.
Exploration of WL loss weight $\lambda_{W\!L}$. To ensure structural symmetry in WF-VAE, we introduce WL loss. As shown in Fig. 7(b), the model substantially decreases PSNR performance when $\lambda_{W\!L}=0$. Our experiments demonstrate optimal results for both PSNR and LPIPS metrics when $\lambda_{W\!L}=0.1$. Furthermore, our analysis of different loss functions reveals that utilizing L1 loss produces superior results compared to L2 loss.
@@
Ablation studies of model architecture. First, We evaluate the effectiveness of the energy pathway by examining its impact when removed from layer 3 alone and both layers 2 and 3. Specifically, we eliminate the wavelet transform layer and inflow-outflow module, setting the concat input to a zero tensor. This analysis highlights the advantages of integrating low-frequency video energy into the backbone. Second, we investigate the importance of the proposed WL Loss in regularizing the encoder-decoder. Third, we analyze the effect of replacing the normalization method with layer normalization [1] for Causal Cache. The results of these ablation studies are shown in Tab. 4.
@@ 4.4 Causal Cache
To validate the lossless inference capability of Causal Cache, we compare our approach with existing block-wise inference methods implemented in several open-source LVDMs. OD-VAE [6] and Allegro [50] offer spatio-temporal tiling inference implementations, while CogVideoX [40] adopts a temporal caching strategy. As demonstrated in Tab. 5, both tiling strategies and conventional caching methods exhibited performance degradation, while Causal Cache achieves lossless inference with performance metrics identical to direct inference.
@@ 5 Conclusion
Acknowledgments
We thank all the anonymous reviewers for their constructive comments. This work was supported in part by the Natural Science Foundation of China (No. 62202014, 62332002, 62425101, 62088102).